KylieAurealis:

Can someone help me with Algebra 2?

Shadow:

Sure. Just post the problems that you need help with.

KylieAurealis:

Number 18

Shadow:

So you need to write an equation for the line shown in the graph. Are you familiar with slope intercept form (shown below) ? y = mx + b

KylieAurealis:

Yes

Shadow:

What does m and b stand for?

KylieAurealis:

M is slope and B is the y intercept

Shadow:

What is the y intercept?

KylieAurealis:

On the problem?

Shadow:

In general, like what is it?

KylieAurealis:

The point where the line crosses the y axis

Shadow:

Correct. So simply locate that point on the graph and you have your b value.

Shadow:

I just wanted to make sure you knew what it was.

KylieAurealis:

So would it be -2?

Shadow:

Correct.

KylieAurealis:

What about m?

Shadow:

m stands for the slope of the line. Do you know what slope is?

KylieAurealis:

The measure of of a line?

Shadow:

Close. It is the rise/run of a line. Let me give you an example. Say I have some stairs.  Each step is three inches long and one inch high. If I were to put a straight line over the top of each step, the slope of it would be 1/3. That's because it rose by one and ran by three.

Shadow:

Note that rise relates to the y axis (vertical) and run relates to the x axis (horizontal). Try to get the imagery in your mind of it rising, then running. That helps you envision the slope well.

KylieAurealis:

So is rise over run? so for a examply 5 would be rise and 2 would be run?

Shadow:

For problem 18?

KylieAurealis:

No just a example. But for 18 would it be m=-2/-5x - 2?

Shadow:

Your rise is correct but your run is not. Try take a look at which direction it's going in.

Shadow:

Also, if a slope looks like this,  Then it is a negative slope. You obtain your slope by dividing two numbers, the rise and the run. If a slope is negative, then only one number (the rise or run) can be negative.

KylieAurealis:

So 5?

Shadow:

Yes. It is moving towards the positive side of the x axis so the five would be positive.

KylieAurealis:

m=-2/5?

Shadow:

Yes.

KylieAurealis:

So y=-2/5 - 2?

Shadow:

Wait I think I'm blind. Lol. http://prntscr.com/klg3p9 Look at these two points. How much is the rise and how much is the run. It's not -2/5. I don't know what points we were looking at, haha.

KylieAurealis:

:/ this is confusing

Shadow:

http://prntscr.com/klg47m This may help. How much did it rise by, and how much did it run by? It may also help to think of it as a triangle.

KylieAurealis:

I cant see it

Shadow:

Look at your graph. Just find two points on the line, then measure by how much it rose, then by how much it ran.

Shadow:

1 attachment
Shadow:

Can you view that?

KylieAurealis:

So it rose one and ran 3?

Shadow:

If it rose by one, it would be on the positive side of the x axis (above it). Instead, where did it go?

KylieAurealis:

It went negative

Shadow:

So what sign would the one have?

KylieAurealis:

-1/3?

Shadow:

Correct

KylieAurealis:

So it would be y=-1/3x - 2?

Shadow:

Yes

KylieAurealis:

Ok thank you!

Shadow:

No problem. Glad you got it.